The mean and standard deviation of a sample of 20 observations were found to be 75 and 2.5 respectively. On checking the original the original figures, it was discovered that one observation which was actually 68, was copied down as 86. Find the correct mean and standard deviation.

Answer by stanbon(75887)

The mean and standard deviation of a sample of 20 observations were found to be 75 and 2.5 respectively. On checking the original the original figures, it was discovered that one observation which was actually 68, was copied down as 86. Find the correct mean and standard deviation.
----
Old mean:: (sum of observations)/20 = 75
sum of observations = 20*75 = 1500
-----------
New mean = (sum of obsers - (86-68))/20 = (1500-18)/20 = 1482/20 = 74.1
-------------
Cheers,
Stan H.
--------------